The life time of a particular brand smartphone's battery is known to have Gamma distribution with αα = 4 and ββ = 2. If we take a random sample of 50 these batteries, what is the probability that the average life time of this random sample will be between 7 and 9 months?
Answer #1
mean μ = αβ = 8
varaince =σ2= αβ2 = 16.00
standard deviation σ=β*√α= 4.0000

since n>=30, we can use normal approximation from central limit theorem:

for normal distribution z score =(X-μ)/σx
here mean=       μ= 8
std deviation   =σ= 4.000
sample size       =n= 50
std error=σ=σ/√n= 0.56569

  probability that the average life time of this random sample will be between 7 and 9 months:

probability =P(7<X<9)=P((7-8)/0.566)<Z<(9-8)/0.566)=P(-1.77<Z<1.77)=0.9616-0.0384=0.9232